OpenStudy (ikou):

Can someone help me? I need to prove this Identity is true (x - y) (y - x) = (xy – x^2 - y^2 - yx)

OpenStudy (whpalmer4):

Just multiply it out... \[(x-y)(y-x) = x(y-x) -y(y-x)\]by distributive property Now distribute again on that and you should be done!

OpenStudy (ikou):

Oh, well any way I could plug numbers in? Like to get 35=35

OpenStudy (ikou):

Somebody please help me, this is taking much longer than it should

OpenStudy (usukidoll):

no need... just expand the (x-y)(y-x) with the distributive property and it's there.

OpenStudy (usukidoll):

start with the left side of the equation.

OpenStudy (usukidoll):

it's like either use the distributive property on the left to get the right hand side of the equation or factor the right hand side of the equation to get the left side of the equation

OpenStudy (usukidoll):

ugh man I fell for that.. nice @whpalmer4 I saw what you did there :/ you used factor by grouping dude. >_< good one :)

OpenStudy (ikou):

Thank you for explaining, but if you, say, make x=# and y=#, what should I do to make them equal to eachother

OpenStudy (usukidoll):

hmmm we could try letting x and y be the same value

OpenStudy (usukidoll):

though I'm trying to do one in my head ... letting x = 1 and y = 1 the result would become (1-1)(1-1)=((1)(1)-1^2-1^2-(1)(1)) (0)(0)=(1-1-1-1) 0=(0-1-1) 0=(-1-1) and that's a false statement

OpenStudy (ikou):

I don't think that would work, because on the left it would be zero right off the bat

OpenStudy (usukidoll):

0 doesn't equal to -2. I know that.. so x and y needs to be different

OpenStudy (ikou):

in looking for something that gives the same number on the left to the right

OpenStudy (ikou):

I've been working on this for days and its really confusing

OpenStudy (usukidoll):

maybe x = 2 , and y = 3 ? (2-3)(3-2)= ((2)(3)-2^2-3^2-(3)(2)) (-1)(0) = (6-4-9-6) 0=(2-9-6) nooooooope I'm assuming that there is no x,y pair and we may need to do with what @whpalmer4 suggested.

OpenStudy (zarkon):

\[(x - y) (y - x) = (xy – x^2 - y^2 \color{red}{+} yx)\]

OpenStudy (ikou):

@Zarkon Why would there be a +?

OpenStudy (zarkon):

\[(-y)(-x)=yx\]

OpenStudy (usukidoll):

DOH! Sign error that makes a big difference.

OpenStudy (usukidoll):

-___________________-!

OpenStudy (usukidoll):

a negative times a negative is positive . Now let's see if we can find an x,y pair that makes the equation equal

OpenStudy (ikou):

If this is what was hanging me up for days I will go into a damn screaming fit

OpenStudy (zarkon):

also, you cant prove it with examples. you need to multiply it out like @whpalmer4 was doing

OpenStudy (usukidoll):

which is either start from the left to get the result on the right hand side of the equation or vice versa right @Zarkon ?

OpenStudy (usukidoll):

(1 - 1) (1 - 1) = (1 – 1^2 - 1^2 +1) (0)(0) = (1-1-1+1) (0) = (0-1+1) 0=(-1+1) 0=0 ok I'm gonna be guilty for doing this because a proof shouldn't be plug x and y but I let x = 1 and y =1 and the equation holds true. It's the sign error that threw everything out of the mix.

OpenStudy (ikou):

Could this come out the same if it was a bigger number like 3?

OpenStudy (whpalmer4):

Yeah, those pesky - signs are the work of the devil! I'm told that in days of old, mathematicians would often write their equations so they didn't need to use them! \[x^2 -3x + 6 = 0\]would be written \[x^2 + 6 = 3x\]and so on.

OpenStudy (usukidoll):

ALso I didn't eat dinner yet. Can't do math without energy. yeah it should be the same

OpenStudy (ikou):

heh, Im so happy this is finally put to rest. Sorry for my sign mistake, this damn thing kept me up till 3 am each night

OpenStudy (whpalmer4):

Actually, I think the original problem is incorrectly stated. \[(x-y)(y - x) = x(y -x) - y(y - x) = x(y-x) - y(-1)(x-y) = x(y-x) + y(x-y) \]\[= xy - x^2 + yx - y^2 = 2xy - x^2 - y^2\]

OpenStudy (ikou):

Why have you dragged me back to hell

OpenStudy (whpalmer4):

\[y-x = -1(x-y)\]so we just have \[-1(x-y)(x-y) = -1(x-y)^2 = -1(x^2 - 2xy +y^2) = -x^2 + 2xy- y^2\] Hey, who copied the problem?

OpenStudy (usukidoll):

wasn't me.

OpenStudy (ikou):

This is for an assignment where I have to create a new identity for polynomials and have to go through proving it

OpenStudy (whpalmer4):

Possibly the problem was to prove that \[(x-y)(y+x)= (xy - x^2 - y^2 - yx)\]? Notice that the right side simplifies to just \(-x^2 -y^2\) because \(xy - yx = xy - xy = 0\)

OpenStudy (usukidoll):

omg I see it now. -_-

OpenStudy (usukidoll):

OMG! THIS IS TROLL LEVEL >:O

OpenStudy (usukidoll):

like xy and yx cancel out

OpenStudy (whpalmer4):

Well, anyhow, to prove an identity, just manipulate both sides of the equation as necessary to meet in the middle with the same thing on both sides...

OpenStudy (ikou):

I need a result where (x - y) (y - x) = (xy – x^ - y^2 + yx) have a result of (example) 20=20

OpenStudy (usukidoll):

leaving (x-y)(y+x) = -x^2-y^2 and then (x-y)(y+x) = -(x+y)(x-y) ............ x+X

OpenStudy (whpalmer4):

Well, you don't prove it works for all numbers by plugging in numbers, though you can prove that it DOESN'T work for all numbers by plugging in numbers.

OpenStudy (ikou):

This assignment is going to be the death of my grade in this class. I really should know how to do this, it really shouldn't be this hard, and theres no + in (y - x)

OpenStudy (whpalmer4):

Well, I'm sorry, but the identity you posted simply is not an identity. Try plugging in \(x =1, y = 2\)

OpenStudy (ikou):

wait a sec, I worked it out and it was -1=-1

OpenStudy (whpalmer4):

\[(x-y)(y-x) = xy - x^2 - y^2 - yx\]is your statement, yes? \[(1-2)(2-1) = (1)(2) - (1)^2 - (2)^2 - (2)(1)\]\[-1*2 = 2-1-4-2\]\[-2=-5\]not in my universe :-)

OpenStudy (ikou):

no it was + yx I screwed up in my signs and (2-1)=1 so -1*1=-1

OpenStudy (whpalmer4):

so for the last time, what is the identity you are trying to prove?

OpenStudy (ikou):

(x - y) (y - x) it's a real pain in the retriceand confusing

OpenStudy (whpalmer4):

no, what is the entire statement? You have no = sign there

OpenStudy (ikou):

ah, sorry im a bit tired, its is (x - y) (y - x) = (xy – x^ - y^2 + yx)

OpenStudy (whpalmer4):

Okay, that's what I was saying earlier:\[(x-y)(y-x) = 2xy - x^2 - y^2\]your version just hasn't collected the \(xy\) terms together Here is a proof: \[(x-y)(y-x) = (xy - x^2 - y^2 + yx)\]we expand the left side twice with the distributive property: \[x(y-x) - y(y-x) = xy - x^2 - y^2 + yx\]\[xy - x^2 - y^2 -(-yx) = xy - x^2 - y^2 +yx\] \[xy - x^2 - y^2 + xy = xy - x^2 - y^2 + yx\]now we change the order of the product for the last term on the left hand side because multiplication is commutative: \[xy - x^2 - y^2 + yx = xy - x^2 - y^2 + yx\] Done.

OpenStudy (whpalmer4):

actually, didn't even need that last step if I hadn't changed \(-(-yx)\) to \(+xy\) by habit of always going in ascending lexicographic order

OpenStudy (ikou):

Thank you for clearing things up! I love you all for helping ♥

OpenStudy (ikou):

Thank you @whpalmer4 senpai!

OpenStudy (whpalmer4):

You're welcome!